### v1

Abstract
We present analyses of the early data from Rubin Observatory’s Data Preview 1 (DP1) for the globular cluster 47 Tuc field. The DP1 dataset for 47 Tuc includes four nights of observations from the Rubin Commissioning Camera (LSSTComCam), covering multiple bands (). We address challenges of crowding near the cluster core and toward the SMC in DP1, and demonstrate improved star-galaxy separation by fitting fifth-degree polynomials to the stellar loci in color-color diagrams and applying multi-dimensional sigma clipping. We compile a catalog of 3,576 probable 47 Tuc member stars selected via a combination of isochrone, Gaia proper-motion, and color-color space matched filtering. We explore the sources of photometric scatter in the 47 Tuc color-color sequence, evaluating contributions from various potential sources, including differential extinction within the cluster. Finally, we recover five known variable stars, including three RR Lyrae and two eclipsing binaries. Although the DP1 lightcurves have sparse temporal sampling, they appear to follow the patterns of densely-sampled literature lightcurves well. Despite some data limitations for crowded-field stellar analysis, DP1 demonstrates the promising scientific potential for future LSST data releases.
1 Introduction
The Vera C. Rubin Observatory, a joint NSF and DOE initiative situated on Cerro Pachón in northern Chile, is designed to carry out the Legacy Survey of Space and Time (LSST; Ivezić et al., 2019). The LSST is a decade-long program capturing rapid, deep, multi-band imaging of the entire southern sky at high spatial and temporal resolution across the six bands. Before full survey operations begin, commissioning efforts produce small, compared to the future official data releases, volumes of good quality data intended to test and refine the observatory’s systems. To support early scientific investigations and community readiness, Rubin offers Data Previews through its Early Science Program (Guy et al., 2025), delivering preliminary but science-grade products and early access to data services.
As the earliest on-sky campaign, the Rubin Commissioning Camera (LSSTComCam) was used to align and verify the end-to-end functionality of the Rubin Observatory’s systems, resulting in 1,792 science-grade exposures collected in late 2024. These data were processed with the LSST Science Pipelines (Bosch et al., 2019; Developers, 2025) to produce Data Preview 1 (DP1)111https://dp1.lsst.io. The available DP1 data products comprise raw and calibrated single-epoch exposures, coadds, difference images, and various detection catalogs, totaling 3.5 TB. It contains 2.3 million distinct astronomical objects, and is accessible via the Rubin Science Platform, a cloud-based data science environment for data analysis. One of the DP1 target fields is centered on the globular cluster 47 Tucanae (47 Tuc).
47 Tuc, aka NGC 104, is the second-brightest globular cluster in the sky, with mag. It’s distance was determined to be kpc using Gaia DR2 parallax by Chen et al. (2018), which corresponds to a true distance modulus of . The line-of-sight reddening toward this relatively high-latitude cluster was estimated at by Hesser & Philip (1976) based on Strömgren photometry of distant early-type stars along that sightline. This low value, consistent with its Galactic latitude of , has been propitious, making it possible for 47 Tuc to be used as a template for stellar populations in a metal rich globular cluster, whereas other metal rich ones are to be found in the Galactic plane and plagued with high reddening that varies on the scale of a cluster.
Its Gaia distance is in agreement with the distance modulus of , derived from model analysis of a detached binary (V69) star in the cluster by Brogaard et al. (2017). Other prior values, such as from the modeling of color-magnitude diagram features, and the cooling curves of its white dwarfs are dependent on knowing or assuming the detailed metal content and the age of 47 Tuc’s stars. Thus Gaia’s parallax based distance sets the stage for a more sure-footed determination of the properties of the cluster’s stellar content, its age, and its history. Despite being an early known metal rich stellar system (e.g., (Rennó et al., 2020) or (Simunovic et al., 2023)), the morphology of its color-magnitude diagram implies an age of between and Gyrs (more specifically Thompson et al. (2020) claim Gyr), making it as old (to within discernible uncertainties) as the oldest globular clusters of the Galaxy. It has been studied using so called “chromosome mapping” of detailed metallicities, which show that its oldest stars likely followed multiple evolutionary and chemical enrichment pathways (Marino et al., 2023). Given its proximity, it also has been a target for studying the cooling sequence of its white dwarfs. Since 47 Tuc is a well-studied cluster, the parameter values listed here are illustrative examples rather than exhaustive or unique measurements.
On the sky, 47 Tuc sits in front of the western extension of the Small Magellanic Cloud (SMC). A concentration of stars situated 15 from the center of 47 Tuc was first reported by Bellazzini et al. (2005) and surmised to be a background star cluster (Bologna A) and later confirmed to be an SMC cluster (Bellazzini et al., 2019). 47 Tuc itself has a highly concentrated core (Shapley class III), with claims that it contains an intermediate mass black hole (Della Croce, A. et al., 2024). By contrast, it also exhibits an extended periphery of stars: e.g. Piatti (2017) has claimed to find stars out to 5.5 times the tidal radius of , that indicate a halo like structure as seen in NGC 1851 (Olszewski et al., 2009). There is a radial gradient in the stellar population (Li et al., 2014), involving metallicities, helium abundance, and mass segregation. These characteristics collectively feed the speculation that like Cen and M54, 47 Tuc is the surviving central remnant from a dwarf galaxy merger.
The complete 6-dimensional phase-space characterization of this globular cluster from Gaia provides details for good dynamical models of its motion through the Galaxy. If we are to use 47 Tuc as a laboratory for studying the detailed processes of its interactions within the Galactic field, it is necessary to trace the stars out as far as possible, but where stars belonging or originating in 47 Tuc must be tagged as distinct from the background of the SMC background stars; those from the ambient “stellar field” of the Galaxy through which 47 Tuc is passing; as well as the stars that may have been tidally stripped earlier from a possible dwarf galaxy progenitor of 47 Tuc. This study is an early look at real data from the Rubin Observatory systems which can inform us of the strengths and limitations of the data, as well as the state of our currently available methodology for attaining the required precision and accuracy in photometry and proper motions for disambiguating the stellar population components.
This paper is organized as follows. In Section 2.2, we briefly describe the observation and the DP1 data products for the 47 Tuc field, and then present the field-specific data properties. Section 3 describes the 47 Tuc member star selection process, explores the photometry of selected member stars, and investigates the LSSTComCam detection of known variable stars. We summarize our conclusions and findings in Section 4.
| Property | ||||||
| aaNumbers in parentheses indicate the subset of visits included in the DP1 final processing. | 6 (0) | 10 (10) | 33 (32) | 19 (19) | … | 5 (5) |
| Observing Nights | … | 2024-12-01 2024-12-07 | 2024-11-25 2024-11-26 2024-12-01 2024-12-07 | 2024-11-25 2024-11-26 2024-12-07 | … | 2024-11-26 |
| bbNumber of CCDs contributing to CCD-level statistics per filter. | … | 90 | 288 | 171 | … | 45 |
| Median Seeing (′′) | … | 1.27 | 1.25 | 1.11 | … | 1.33 |
| Median Limiting Mag | … | 24.14 | 24.14 | 23.77 | … | 21.66 |
| Maximum Limiting Mag | … | 26.15 | 27.07 | 28.63 | … | 23.03 |
| Areal Coverage (deg2)ccNumbers in parentheses indicate the areal coverage with the central hole excluded. | … | 0.980 (0.825) | 1.064 (0.915) | 1.166 (0.997) | … | 0.807 (0.747) |
2 Data for the 47 Tuc Field
2.1 Observations
The 47 Tuc field was one of seven target fields observed during the LSSTComCam (the Rubin Observatory Commissioning Camera; SLAC National Accelerator Laboratory & NSF-DOE Vera C. Rubin Observatory, 2024) on-sky campaign (24 October–11 December 2024), and is centered at (RA, Dec) = (6.128, –72.090) degrees. Observations were obtained in five bands (), with the deepest coverage in the band. Imaging employed random dithers within a 02 radius and occasional rotator offsets. The dense stellar environment of 47 Tuc makes it a valuable testbed for wavefront sensing, internal calibrations, and crowded-field stellar population studies. Although saturation in the cluster core led to missing coadds and prevented reliable measurements in that region, this target field still offers a great dataset for scientific analysis. The LSSTComCam observations for the 47 Tuc field include 6 visits in the band, 10 in , 33 in , 19 in , and 5 in . However, all -band visits and one -band visit failed to meet the quality criteria and were excluded from the final DP1 processing. Details about the LSSTComCam on-sky campaign and the DP1 can be found in SITCOM-149 (On behalf of Rubin Observatory Project, 2025) and DP1 paper (NSF-DOE Vera C. Rubin Observatory, 2025a).
In Figure 1, we present the distributions of the delivered image quality and depth across individual CCD images for the , , , and bands. The left panel shows the distribution of PSF FWHM, with most images having seeing between 0.8″and 2.0″, and a peak around 1.1-1.3″across all bands. The right panel shows the distribution of 5 limiting magnitudes, with the median depths reaching approximately 24.14, 24.14, 23.77, and 21.66 mag in the , , , and bands, respectively. A per-CCD summary of observational properties by filter is listed in Table 1.
2.2 DP1 Data Products for 47 Tuc
All data were processed with the LSST Science Pipelines software (Bosch et al., 2019; Developers, 2025) at the U.S. Data Facility (USDF; housed at the SLAC National Laboratory). Initial processing steps in instrument signature removal (ISR) include applying bias, dark, and flat-field frames, corrections for crosstalk and the brighter-fatter effect (Broughton et al., 2024), and removal of other instrumental artifacts and effects (see Plazas Malagón et al. 2025 for details about the ISR procedures, and Fagrelius & Rykoff 2025 for more information on the calibration data products). Source detection and deblending is then run on the ISR-corrected images, followed by measurements of source centroids, shapes, and photometry. Initial photometry and shape estimates are used for the selection of stars for point-spread function (PSF) fitting; a preliminary fit to the PSF is then determined using a modified version of PSFEx (Bertin, 2011), after which a second, more careful and accurate, round of PSF fitting is performed using a modified version of Piff (Jarvis et al., 2021a, b). Initial photometric and astrometric calibrations were performed by cross-matching to the Monster reference catalog – an amalgam based on Gaia DR3 astrometry and photometry, and also including transformed DES, PanSTARRS1, SkyMapper, VST, and SDSS photometry (Ferguson et al., 2025). Subsequently, more detailed determination of the astrometric solution uses fitting routines from gbdes (Bernstein et al., 2017; Bernstein, 2022), and a global photometric calibration is performed using the Forward Global Calibration Method (FGCM; Burke et al. 2018). Additional steps such as background estimation and subtraction are performed, and detection and measurement steps are repeated on the final per-detector visit_images produced by the single-visit processing.
In full data release processing, all visit_image datasets are resampled onto a common sky projection (defined by a skymap), convolved to a common PSF, and then stacked (with inverse-variance weighting) into a “coadd” image. Outlier rejection is performed to prevent image artifacts, cosmic rays, and moving objects from printing through to the final coadd images. Object catalogs (NSF-DOE Vera C. Rubin Observatory, 2025b) are produced from detection and measurement of objects on the final deep_coadd images (NSF-DOE Vera C. Rubin Observatory, 2025c).
Details of the data products released with Data Preview 1 (DP1; NSF-DOE Vera C. Rubin Observatory, 2025d) are provided in https://dp1.lsst.io. Briefly, the DP1 data products include raw images, processed per-detector single-frame visit_images and their corresponding Source catalogs, and stacked deep_coadd images and their corresponding Object catalogs, among others.
2.2.1 Depth and Coverage
Figure 2 shows the cumulative -band imaging depth (i.e., coadd depth), expressed as the 5 limiting magnitude, for the 47 Tuc field. This map effectively represents an idealized PSF magnitude limit, assuming perfect detection and measurement. The outlines of coadd patches are overlaid, with each tract distinguished by a unique color and line style. The bright core of 47 Tuc is saturated in the LSSTComCam commissioning data, leading to a lack of coadd coverage in that region. The absence of coadds in that region implies that no object measurements, difference imaging, difference-image source analysis, and forced photometry analysis could be performed.
To estimate the per-band areal coverage of the coadd images of the field, we identified valid regions in the depth map and traced their boundaries using a contour-based approach. This method captures the detailed shape of the coverage, including any internal gaps or holes. The resulting contours were converted to sky coordinates and used to construct a polygonal representation of the footprint. The total area was then computed from this geometry both including and excluding the saturated central region. The outermost boundary and the central hole determined for the band are shown in Figure 2. The outer boundary traces the full extent of the valid data, while the inner boundary delineates the saturated region excluded from the hole-corrected area measurement. The resulting sky coverage in the band is approximately 1.064 deg2 with the central hole included, and 0.915 deg2 when the hole is excluded. The areal coverage in other bands is listed in Table 1.
2.2.2 Impact of Crowding on Deblending
In this and following sections, we explore the Object table to provide an overview of selected properties that are most relevant for analyzing the stellar populations in this crowded field. The Object table stores measurements of objects detected in the coadd images with a signal-to-noise ratio (SNR) of 5 in at least one band, and contains 1,296 columns (NSF-DOE Vera C. Rubin Observatory, 2025b).
The ADQL query used in this analysis is as follows:
The query returned 115,493 objects.
Figure 3 shows the i-band 5 limiting magnitude map (left), the number count map of -band measured objects in the Object table (middle), and the number count map of all Gaia DR3 objects within 1 degree from the 47 Tuc field (right). As mentioned earlier, the depth map is estimated from the properties of the images contributing to the coadds assuming the ideal detection and measurement scenario, while the density map include all entries in the Object table for detections in the coadd images with SNR 5 in at least one band. In contrast to the limiting magnitude map, the Object table density map is discontinuous, with noticeable gaps in addition to the expected central hole, and a sharp decline in the number of detected objects near the field center. The decline becomes more pronounced toward increased RA, consistent with enhanced stellar contamination from the SMC in that region, as shown in the Gaia density map. The SMC is centered at (RA, Dec) = (13.187, -72.8297) degrees, with the angular separation of 2.578 degrees from the 47 Tuc field.
To understand the origin of the apparent lack of sources in the central and eastern parts of the field, we examine deblending outcomes for two representative tract/patch pairs: (454, 48), located on the eastern side where the resulting object density is low, and (453, 35), on the western side, which includes both dense and sparse regions. Figure 4 shows the spatial distribution of sources for which the deblending process was either performed or skipped due to excessive memory consumption or runtime requirements. In the eastern region (left panel), the notable drop in Object table source counts appears to be primarily driven by skipped deblending, likely due to crowding associated with the proximity to the SMC. Deblending can be skipped for various reasons, including excessively large parent footprints, too many detected peaks, a single-peak configuration, or excessive masking. In the western region (right panel), deblending was largely skipped in the crowded inner area but completed successfully in the outer parts, where the intrinsic object density is lower. This underscores the inherent challenges of performing photometry in crowded fields, a regime for which the LSST Science Pipelines are not specifically optimized. Improvements to scarlet (Melchior et al., 2018) lite should allow processing to push further into these and other dense regions before the first data release.
2.2.3 Star-Galaxy Separation
Unresolved and resolved objects often occupy the same part of the sky but represent fundamentally different physical systems. Misclassifying galaxies as stars (or vice versa) leads to systematic errors in object catalogs and thus biased measurements of stellar populations, galaxy properties, and large-scale structure. In crowded fields like 47 Tuc, where stars and compact galaxies overlap in morphology and color especially at fainter magnitude, robust classification becomes especially critical for downstream science.
The LSST Science Pipelines report an extendedness parameter, which is based on the ratio of cModel and PSF magnitudes; point sources should have roughly equal fluxes from these two measures, while extended sources will have more flux in the model measurement than the PSF. Objects are classified as “not extended” (extendedness=0) if the ratio between their PSF and cModel fluxes satisfies . In this work, we use the refExtendedness parameter, which for each object corresponds to the extendedness measured in the “reference band” (selected as described in Sec. 3.4 of Bosch et al. 2018).
Figure 5 shows vs. color-color diagrams (top panel) and vs. color-magnitude diagrams (bottom panel), separated by extendedness classification. The left column displays likely stars (refExtendedness 0) using PSF magnitudes, while the right column shows likely galaxies (refExtendedness 1) using cModel magnitudes. In the color-magnitude diagram of likely stars, three components are evident: 47 Tuc, the SMC, and the Milky Way. We overlay PARSEC v1.2S isochrones on the stellar locus of 47 Tuc and the SMC (Bressan et al., 2012; Tang et al., 2014; Chen et al., 2015). The red isochrone represents the old stellar population of 47 Tuc, with an age of 12.4 Gyr, [M/H] = -0.5, and distance of 4.66 kpc (Simunovic et al., 2023). The blue isochrone traces diffuse stellar populations from the SMC, with an age of 8 Gyr, [M/H] = -1, and distance of 62 kpc. Traces of both the 47 Tuc and SMC sequences remain visible in the color-magnitude diagram of extended objects, indicating that star-galaxy separation using the refExtendedness parameter becomes unreliable at fainter magnitudes in crowded fields such as 47 Tuc.
Color-color diagrams are a powerful tool for separating stars from unresolved background galaxies, as stars occupy well-defined, continuous loci in multiple color-color spaces, while galaxies and other contaminants typically deviate from these loci. This distinction becomes especially useful in deep imaging where morphological classification becomes unreliable at faint magnitudes. To leverage this, we performed polynomial fits to the stellar locus in three color–color spaces: , , and .
Figure 6 illustrates the process and compares the multiple color-color diagram-based star–galaxy separation against that based on the refExtendedness parameter. The band was excluded due to its limited depth. For each pair, we selected likely 47 Tuc stars using the isochrone presented in Figure 5 (see Section 3.1.1) and moderate color ranges to exclude extreme outliers and poorly measured objects. A fifth-degree polynomial was fit to each observed stellar locus. We then computed the residuals of all objects from the fitted curve. Objects with residuals exceeding 5 were identified in each color–color space, allowing us to exclude sources that lie significantly off the stellar locus in all three spaces. The resulting polynomial fits are:
In the bottom panels of Figure 6, we compare the star/galaxy separation using the pipeline-measured refExtendedness parameter and our multi-dimensional sigma-clipping method. The refExtendedness-based classification misidentifies highly-likely, relatively bright stars along the isochrone as extended sources, leaving an imprint of the 47 Tuc stellar locus in the extended-object color-magnitude diagram, as shown in the bottom right panel of Figure 5. In contrast, our approach more reliably excludes faint objects with colors inconsistent with the stellar locus, likely corresponding to background galaxies.
3 Analysis
3.1 Selecting 47 Tuc Member Stars
3.1.1 Isochrone-based Selection
As shown in Figure 5, the 47 Tuc sequence is well defined and closely follows the model isochrone in the color-magnitude diagram. We therefore identified likely member stars by measuring their proximity to a reference isochrone in color-magnitude diagram space. For each observed star, we computed the minimum distance to the isochrone, normalized by uncertainties in both color and magnitude. Tighter color tolerances (0.05 mag) were applied to bright stars ( 20 mag) to reflect their higher photometric precision, while fainter stars used a broader threshold of 0.1 mag.
We converted these distances into a likelihood-like score using a standard normal distribution. The closer the star is to the fiducial isochrone, the higher the matched scores. To identify probable member stars, we selected those exceeding threshold of 0.2. This method effectively isolated high-confidence candidates in the color-magnitude diagram space, leaving 4,276 candidate stars. However, in regions of the color-magnitude diagram where 47 Tuc main sequence overlaps with the SMC red clump, this approach alone is insufficient. To mitigate contamination in that regime, we applied an additional cut based on Gaia DR3 proper motion measurements (Gaia Collaboration et al., 2023).
3.1.2 Gaia Proper Motion-based Selection
We retrieved Gaia DR3 proper motions and photometry using the same center and radius as the Object table query, excluding sources with extreme proper motions or astrometric_excess_noise 0.2, yielding 102,902 objects. Given Gaia’s limiting magnitude of , the goal in this step is not to identify stars consistent with the 47 Tuc proper motion, but to exclude clear outliers in the proper motion space from the isochrone-selected sample.
Figure 7 demonstrates the separation of stellar populations in proper motion space based on a Gaussian Mixture Model (GMM) classification. We applied the GMM to the distribution of proper motions in RA () to identify kinematic components in the 47 Tuc field using sklearn.mixture.GaussianMixture. The model was fit using only the values, assuming three Gaussian components to represent the Milky Way contaminants, the 47 Tuc cluster, and the background SMC populations. The resulting classification was then used to label stars and examine the corresponding distributions in DEC proper motion (). Although the GMM was trained solely on , the distributions for each component remain well distinct, further validating the classification. In the top panels, we present these distributions by overplotting scaled Gaussian kernel density estimates of and for each component identified by the -based GMM, weighted by the relative number of stars assigned to each group. The bottom panels show spatial density maps of Gaia DR3 objects with proper motions consistent with the 47 Tuc component (left) and the SMC component (right). A total of 308 stars were excluded from the isochrone-selected sample during this proper motion filtering step.
3.1.3 Selection in Color-color Diagrams
To further refine the membership sample, we applied the same sigma-clipping method in the three color-color spaces, using the polynomial fits determined in Section 2.2.3. This multi-dimensional color-color filtering removed a total of 392 objects that deviate significantly from the stellar locus, helping eliminate any remaining contaminants with inconsistent colors.
Figure 8 presents a summary of the sequential member selection process. The first three panels show color-magnitude diagrams illustrating the impact of each filtering step, from left to right: isochrone-based pre-selection, proper motion refinement, and final color-color diagram filtering. Each panel highlights the progressive narrowing of the stellar locus and removal of contaminants. The fourth panel compares the -band magnitude distribution of the initial isochrone-selected sample and the final sample after all filtering steps. The proper motion refinement effectively removed significant contamination from the SMC, particularly around its red clump, as well as brighter stars down to Gaia’s depth limit. In contrast, the color-color diagram filtering step primarily excluded faint objects—likely background galaxies—that could not be identified through Gaia proper motion due to their low brightness. The final sample consists of 3,576 member stars. The list of the selected 47 Tuc member stars is available online.
3.2 Exploring Photometry of 47 Tuc
3.2.1 Saturation Limit
Saturation limits define the upper reliability threshold of the data and are essential for quality control, completeness assessment, and robust scientific interpretation. Saturated stars yield inaccurate or unusable flux measurements, and without a clear understanding of the saturation limit, bright sources can be mischaracterized, biasing downstream analyses. Even with identical exposure times (as in the DP1 case) and identical spectral types, the saturation limit of a star can vary with seeing, total filter throughput, background level, and other observational conditions. Moreover, as discussed in Section 2.2.2, the LSST Science Pipelines in DP1 were not particularly optimized for crowded fields, unfortunately leading to skipping the deblending process for sources even under modest crowding. Consequently, the brightest star with reliable measurements in the 47 Tuc field does not necessarily reflect the true saturation limit.
The brightest 47 Tuc member star (objectId: 579576081061783277) is on the red giant branch phase and has magnitudes of , , , and in the , , , and bands, respectively. The star has a Gaia counterpart, and its transformed photometry from the Gaia and to the LSSTComCam system (Tucker, 2025) agrees well within a few hundredth of a magnitude in the , , and , where no visit images have saturated pixels. In the band, however, the transformed magnitude differs by 0.06 mag, coinciding with mild saturation in the central pixels of a few visits. This offset is likely due to the fact that the band has the best seeing in the field (see Figure 1 and Table 1), leading to saturation in the core, and the RGB star peaks in flux near that band. The second brightest -band magnitude without saturated pixels is 15.19420.0002. In short, the unsaturated bright limits reported in the Object table for 47 Tuc are approximately 15.70, 15.08, 15.19, and 14.74 mag in the , , , and bands, respectively.
As a check on the onset of nonlinearity in the photometry at the bright end, we examined the color-magnitude diagrams of the low Galactic latitude field Rubin_SV_095_-25. The sequence of numerous red disk turnoff stars in the color-magnitude diagrams of this field departs from its expected vertical shape at the bright end at magnitudes of 15.9,16.2,15.8,15.7 , an indication of the point at which non-linearity begins to affect the photometry of bright stars. These limits are up to a magnitude fainter than seen in 47 Tuc, such that care should be taken with photometry brighter than 16th magnitude.
3.2.2 Photometric Uncertainty
Figure 9 demonstrates the photometric uncertainty of the selected member stars in each band. While the shallowest -band photometry show relatively large uncertainty, the other three bands achieve 1% level precision roughly up to 21.5, 22, 21.5 mag in the , , bands, respectively. The multiple quantized sequences seen in each band reflect variation in image depth across the field. Photometry from deeper exposures (i.e., more input images) yields smaller uncertainties. To illustrate this, each panel is color-coded by the number of input images contributing to the measurement at each object’s position. The distinct sequences correspond to different numbers of input images, confirming that depth is the dominant factor shaping the structure of these sequences.
3.2.3 Analysis of Photometric Scatter
The simple stellar population of 47 Tuc affords the opportunity to diagnose possible sources of scatter in the DP1 photometry. Figure 10 shows the vs. color-color diagram for bright stars (), which features a tight linear sequence. We considered four possible sources to explain the scatter about this sequence: 1) PSF-fitting errors stemming from known Poisson and other observational noise sources; 2) crowding errors (e.g., Olsen et al., 2003); 3) differential extinction by dust (e.g., Pancino et al., 2024); and 4) unknown sources of observational systematic error.
For this analysis, we used the Padova Observatory CMD 3.8 web interface (https://stev.oapd.inaf.it/cmd) to generate a simulated stellar population with age 12.4 Gyr, [M/H]= (Simunovic et al., 2023), and total mass of 104 . The stellar population was drawn from isochrones based on PARSEC release v1.2S+COLIBRI S_37+S_35+PR16 (Bressan et al., 2012; Chen et al., 2014; Tang et al., 2014; Chen et al., 2015; Rosenfield et al., 2016; Marigo et al., 2017; Pastorelli et al., 2019, 2020) and uses the LSST throughputs Release 1.9 (Sept 2023) to compute simulated magnitudes in the LSST passbands. Figure 11 shows the -band luminosity function of the simulated population compared to the observations, where we have normalized the simulation to match the observed number of stars with and adopted a distance modulus and mag (Simunovic et al., 2023).
Our approach was to add the first three sources of photometric scatter described above (PSF-fitting errors, crowding errors, and differential extinction) to the simulated stellar population, fit a line to the vs. relation, and measure the scatter about this relation in to compare with the observations. To add the PSF-fitting errors, for each simulated star we picked a random star with the -band magnitude within 0.1 mag from the observed catalog, looked up its measured photometric error, and picked a random value from a Gaussian distribution with spread equal to the observed error value.
To add errors due to crowding, we used the equations developed by Olsen et al. (2003) to compute the crowding errors in the and bands for each simulated star. These equations require the full simulated input luminosity functions in the three bands and the median image quality measurements from Table 1. With the crowding error scale computed from these equations as a function of magnitude, we assigned a random error to each simulated star from a Gaussian distribution of the appropriate dispersion. We accounted for the correlated errors in crowding between bands in this calculation, such that the crowding contribution to the scatter in is slightly lower than that expected from the crowding error in individual bands.
To add scatter from differential reddening, we again drew random values from a Gaussian distribution in centered on mag, where we tested various values for the width for this distribution, but truncated the distribution at 0 to avoid negative extinction values. We used = 1.18, 0.87, and 0.66 for and , respectively, based on the Cardelli et al. (1989) extinction law with and the LSST filter transmission curves (Yoachim et al., 2025).
Figure 12 shows the relative contributions to the scatter in of the simulated population about the linear fit in the vs. plane, compared to the observed scatter. As seen in the figure, PSF-fitting errors and crowding errors are insufficient to explain the observed scatter in for stars brighter than . Simulated differential extinction with mag is, however, able to explain the observed scatter. We note that this scale of differential extinction is a factor of two lower than the mag measured by Pancino et al. (2024), based on archival wide-field photometry.
We do not claim to have measured differential extinction in 47 Tuc based on modeling the photometric scatter in DP1, but do conclude that in the areas covered, differential extinction can not be as large as seen in Pancino et al. (2024). Those authors cited camera-specific systematic errors as limiting their measurement, and we cannot rule out such errors in DP1. Figure 13 shows the observed median values of the offsets about the line fit to the vs. plane. The figure shows spatially coherent variations in at scales of deg, which matches the scale of the patches in the DP1 reduction. There is not, however, obvious correlation with patch boundaries; more data will be needed to understand the source of photometric scatter in 47 Tuc.
3.3 Variables in the 47 Tuc Field
As an assessment of the pipeline performance and DP1 data products, we crossmatched the 47 Tuc Object and DiaObject (NSF-DOE Vera C. Rubin Observatory, 2025e) tables to a catalog of known variables from Weldrake et al. (2004). The original catalog contains 100 variables stars across an approximately 1 square degree field centered on the cluster, which is presented in their paper in Table 2. Their Tables 3–5 then detail the fitted period and amplitude of of each variable, broken out by variable type: RR Lyrae (RRLyr), Eclipsing Binaries (EcBs) and others, mostly long period variables (LPVs). Separately, a table was published to SIMBAD containing additional photometry, as well as an identifier column and links to other references for each variable. We consolidated the tables, ignoring rows with only limits or approximations of the period, amplitude or photometry, leaving 72 variables remaining. The 72 variables left include 19 EcBs, 8 detached EcBs, 11 LPVs, and 36 RRLyr. All RRLyr recovered likely belong to the SMC according to distance moduli derived by Weldrake et al. (2004).
We then used the Large Survey Database (Caplar et al., 2025, LSDB) to crossmatch the 72 variable catalog to both the Object table and DiaObject table narrowed down to contain only those located in the 47 Tuc field. We find 5 matches to Objects and 62 matches with DiaObjects. Three variables, all RR Lyrae (ID: V23, V73, and V88) are found in both tables. The two objects not found in the DiaObject table correspond to one EcB and one detached EcB. As discussed in Section 2.2.2, the skipped debelending on the denser areas closer to the cluster core results in fewer detected objects toward the center, as shown in the middle panel of Figure 2. In contrast, more DiaObjects are detected closer to the cluster core because the difference images are less crowded after template subtraction. For our variable star crossmatching, we therefore use the DiaObject table, particularly because we expect there to be residual flux in the difference images due to the variability of our targets.
In Figure 14, we show the science, template, and difference images in a small field within 47 Tuc from visit 2024112400101 and detector 3. The overplotted cyan crosses show the diaSources detected in the difference image. The source marked with the red circle is a known RR Lyrae variable star and shows a clear negative flux in the difference image. It appears some of the diaSources may be incomplete subtractions of the brightest stars. At least one of the detections appears to be on a diffraction spike of the brightest star in the image.
There are several flux measurements available in the DP1 datasets. The scienceFlux from the DiaSource table is the forced PSF flux measured on the direct image at the position of the diaSource. We also have forced photometry measurements at the position of the associated diaObject on the difference image (psfDiffFlux) and the preliminary visit image (psfFlux) available from the ForcedSourceOnDiaObject table.
In Figure 15, we plot one of the RRLyr stars (diaObjectId = 579584602276888841; ID: V23) recovered in our crossmatch. Because the 47 Tuc field was only visited on 4 days (Table 1), the lightcurve is quite sparse. However, the shape of the lightcurve looks like a typical RR Lyrae profile and matches the phase-wrapped lightcurve of the same source from Weldrake et al. (2004). We plotted the forced photometry on the visit images at the diaObject position (psfFlux), as there are more flux measurements in the ForcedSourceOnDiaObject table than in the DiaSource table. This is because the source is not always detected in the difference images, for example when its science flux is close to the flux in the template image, and thus does not always make it into the DiaSource table. The phase-folded lightcurve shows a consistent pattern with the densely-sampled one from Gaia DR3 photometry (Clementini et al., 2023).
4 Summary
This study presents a detailed exploration and constructive assessment of Rubin Observatory DP1 data centered on the globular cluster 47 Tuc. The LSSTComCam imaging provided valuable early photometric measurements, while also revealing challenges from crowding, particularly near the core of 47 Tuc and toward the SMC. Although crowding affected deblending and subsequent measurements, improvements to scarlet lite are anticipated to enhance processing capabilities in crowded regions before the first data release.
We identified 3,576 high-confidence 47 Tuc member stars through a rigorous selection process combining isochrone-based filtering on the (, ) color-magnitude diagram, Gaia DR3 proper-motion filtering, and multi-dimensional color-color filtering. By fitting fifth-degree polynomials to the observed stellar locus in three unique color-color diagrams, we achieved better star-galaxy separation than the provided morphology-based extendedness parameter. We also characterized the saturation limits, with caveats in the DP1 photometry in the 47 Tuc field. Further analysis of photometric scatter suggested possible contributions from differential extinction or observational systematics, indicating areas for future investigation with more data.
Further, we successfully crossmatched known variable stars within the 47 Tuc field against the DP1 data, recovering three RR Lyrae stars and two eclipsing binaries. Despite sparse temporal sampling, forced photometry at their fixed locations in every image aligns pretty well with the denser lightcurves from the literature, underscoring Rubin’s promising capabilities for variability detections even in crowded regions.
Overall, while challenges remain, the DP1 data around 47 Tuc convincingly showcase Rubin Observatory’s strong potential for detailed stellar population and variability analyses in crowded stellar fields. Continued improvements to the Rubin Science Pipelines and in-kind programs dedicated to crowded-field stellar photometry are expected to deliver even higher-quality results in future DP2 and DR1.